OpenStudy (anonymous):

Solve for x. 5x - 11 = 7x - 3

OpenStudy (angela210793):

5x-7x=-3+11 -2x=8 x=-8/2 x=-4

OpenStudy (jamesj):

Seriously, you try and do this yourself. Look at the last one I answered for you and imitate the steps.

OpenStudy (anonymous):

Im sure i COULD do it. But I ask easy Q's for easy A for easy Medals (: So People on lower levels can get higher